2-3-B for GPA You previously calculated a GPA using a loop; in this assignment youll use a for loop A GPA, or Grade Point Average, is calculated by summing the grade points earned in a students courses and then dividing by the total units The grade-points for an individual course are calculated by multiplying the units for that course by the appropriate factor depending upon the grade received A receives 4 grade-points B receives 3 grade-points C receives 2 grade-points D receives 1 grade-point F receives 0 grade-point Your program will have a for loop to calculate multiple GPAs and a for loop to collect individual grades (i.e. a nested for loop) For your demo, calculate the GPA to 2 decimal places for these two course combinations First Case 6 units of A 6 units of B 3 units of C Second Case 4 units of A 6 units of B 5 units of C Can I have the code in Python and the output
Answer #1

#python3 code
#please refer to the screenshots and take care of indentations

n = 2
gmap = { 'A': 4, 'B': 3, 'C': 2, 'D': 1 }

for i in range(0,n):
name = input("Enter the student name: ")
s = 0
total_units = 0
for j in range(0,3):
g = input("Enter a grade: ")
units = int(input("Enter units of grade: "))
j = j+1
s = s + (gmap[g.upper()]*units)
total_units = total_units + units
gpa = s/total_units
print("The GPA for", name, "is ",end="")
print("%.2f" %gpa)
i = i + 1


Terminal (100%) ,) 12:32 PM ubuntu En new.c Skewness.java stats.cpp new.py new.cpp 1 #python3 code #please refer to the scree